The length of one of the equal legs of an isosceles triangle is 8 cm less than 4 times the length of the base. If the perimeter is 29 cm, find the length of one of the equal legs.
I am thinking the formula would at least maybe start out as: 2(4x-8) + x = 29 ???

Answer by ankor@dixie-net.com(12692)

That's right:
:
2(4x-8) + x = 29
:
8x - 16 + x = 29
:
9x = 45
:
x = 5
:
Check 4(5)-8 = 12; 12 + 12 + 5 = 29